A toy of mass 0.155 kg is undergoing simple harmon
Answer #1

part A

the total mechanical energy is equal to

E=U_{k}+K=\frac{1}{2}k\cdot x^{2}+\frac{1}{2}\cdot m\cdot v^{2}  elastic potential energy and kinetic energy

Where:

elastic constant

k=305\frac{N}{m}
elongation

x=1.15\cdot 10^{-2}m
speed

v=0.305\frac{m}{s}
mass

m=0.155kg

evaluate numerically

E=\frac{1}{2}\cdot 305\frac{N}{m}\cdot \left ( 1.15\cdot 10^{-2}m \right )^{2}+\frac{1}{2}\cdot 0.155kg\cdot \left ( 0.305\frac{m}{s} \right )^{2}=0.0274J

E=0.0274J

Part B

The total mechanical energy is equal to

E=\frac{1}{2}\cdot k\cdot A^{2}

where:

A is the amplitude

expression for the amplitude.

E=\frac{1}{2}\cdot k\cdot A^{2}\Rightarrow A=\sqrt{\frac{2\cdot E}{k}}

evaluating numerically

A=\sqrt{\frac{2\cdot 0.0274J}{305\frac{N}{m}}}=1.34\cdot 10^{-2}m

part C

mechanical energy is also equal to

E=\frac{1}{2}\cdot m\cdot v_{max}^{2}

Expression for the maximum speed

E=\frac{1}{2}\cdot m\cdot v_{max}^{2}\Rightarrow v_{max}=\sqrt{\frac{2\cdot E}{m}}

evaluating numerically

v_{max}=\sqrt{\frac{2\cdot E}{m}}=\sqrt{\frac{2\cdot 0.0274J}{0.155kg}}=0.595\frac{m}{s}

finally

v_{max}=0.595\frac{m}{s}